# Eye Gaze as a Signal for Conveying User Attention in Contextual AI Systems

### v1

@@ Abstract.
Advanced multimodal AI agents can now collaborate with users to solve challenges in the world. We explore eye tracking’s role in such interaction to convey a user’s attention relative to the physical environment. We hypothesize that this knowledge improves contextual understanding for AI agents. By observing hours of human-object interactions, we first measure the relationship between an eye tracker’s signal quality and its ability to reliably place gaze on nearby physical objects. We then conduct experiments which relay the user’s scanpath history as additional context querying multimodal agents. Our results show that eye tracking provides high value as a user attention signal and can convey information about the user’s current task and interests to the agent.
@@ 1. Introduction
Artificial intelligence (AI) agents have become more connected with users in daily life (Wienrich and Latoschik, 2021), especially by observing context about the user’s prior actions or current world state (Zhang et al., 2024a). New innovations, such as vision-language models (VLMs) (Li et al., 2024) and persistent interfaces, pave the way towards contextual AI agents which “see” the nearby physical world to better assist users. Current AI agent perception of image / video differs greatly from human understanding, so these models often misinterpret context and can respond inappropriately, conflicting with user intent.
Eye gaze is hypothesized as a valuable signal for conveying intent to agents (Ajanki et al., 2010; Büschel et al., 2018; Burlingham et al., 2024b; Zhang et al., 2024b). Gaze communicates information about objects users are interested in, cognitive load, the task being performed, etc. (Mahanama et al., 2022), which can all improve models’ understanding. While eye tracking (ET) is being commonly used as a cursor in extended reality (XR) systems (Plopski et al., 2022), the use of ET in human-agent interactions has only been lightly explored (Sendhilnathan et al., 2024).
A primary benefit from ET is the ability to convey to an agent what the user is or has been interested in. Eye trackers are limited in their accuracy due to a number of factors (hardware / software, slippage, individual user differences, etc.) (Ehinger et al., 2019), constraining which fixated objects can be reliably identified. Objects with small visual area require tighter gaze estimation, so could be unreliable to identify if ET accuracy is insufficient. This work presents a deeper analysis of the requirements and benefits of ET in wearable contextual AI. Using a dataset of egocentric recordings in natural scenarios, we quantify the expected ET accuracy thresholds for detecting physical objects. We then conduct a number of experiments supplying contextual information from ET to VLM queries, reinforcing ET’s value in this space.
@@ 2.1. Eye tracking for selection in extended reality
In recent years, eye gaze has gained popularity as a key signal for HCI in XR systems (Plopski et al., 2022). Eye gaze has comparable usability to controllers (Luro and Sundstedt, 2019; Zhang et al., 2019; Fernandes et al., 2024) while freeing the hands for other tasks and being preferable to users (Piening et al., 2021). While ET is prone to a midas touch fallacy, where false selections are made during ambient fixations (Jacob, 1995), novel HCI methodologies (Khamis et al., 2018) overcome this and make ET an ideal signal for interface navigation. While ET’s value is proven for explicit selection, our work explores its value in contextual AI agent interaction, particularly the ability to implicitly convey a user’s attention and intent (Sendhilnathan et al., 2024). This concept is only lightly explored with AI agents, though eye gaze has been used to facilitate automatic contextual displays (Toyama et al., 2012).
@@ 2.2. Eye gaze encodes scene understanding
Eye gaze and eye movements indicate a viewer’s internal processing of a scene (Yarbus, 1967), reflecting cognitive state and attentional focus as one interprets new visual stimuli (Langton et al., 2000; Eckstein et al., 2017). The sequence of gaze fixations (i.e., scanpath) encodes contextual cues as to future objects of interest (Itti and Koch, 2001; Burlingham et al., 2024a); a number of works have leveraged scanpath history for short-term gaze prediction / anticipation (Huang et al., 2018; David-John et al., 2021; Hu et al., 2021; D’Amelio et al., 2024). Burlingham et al. found that temporal dependencies in scanpaths last for 4-5 fixations on average, and vary substantially among task contexts (Burlingham et al., 2024b). Contextual AI models may be able to leverage this rich, multiscale structure in scanpaths (analogous to how large language models (LLMs) leverage multiscale structure in language), enabling implicit inferences about user intent.
Insights about cognitive encoding of nearby objects can inform our expectations for eye movements in contextual AI (Tatler et al., 2011). For example, humans would tend to look at a coffee mug just before grasping, as the location and orientation of the handle must be encoded before successfully picking up. Objects are prioritized relative to egocentric positions around the viewer, with greater affordances given to objects which are close and candidates for interaction (Costantini et al., 2010; Tatler et al., 2011). The visual system elicits responses to reachable 3D objects, even when there is no intent to interact (Iachini et al., 2014, 2023). So, by analyzing eye gaze fixations on near-field objects, we are able to register a list of possible interactable objects noticed by the viewer.
@@ 2.3. Eye tracking in contextual AI
Information from the physical world can greatly improve user interactions with AI agents (Zhang et al., 2024a). Emerging products, such as the Ray-Ban Meta [^1] https://www.meta.com/smart-glasses and Google’s Ask Photos [^2] https://blog.google/products/photos/ask-photos-google-io-2024/, use image context to improve user interaction. Eye gaze could be fundamental for narrowing the scope of information presented to contextual agents (Ajanki et al., 2010; Büschel et al., 2018), enhancing contextual understanding and avoiding hallucination (Cui et al., 2023; Leng et al., 2024).
Some wearable contextual AI prototypes have been proposed, using eye gaze in different ways to facilitate better querying (Zhang et al., 2024b). For example, the GazeGPT system projects 2D gaze onto an image capture, cropping image contents before interfacing with a VLM. They show gaze-based querying to be faster, more accurate, and more natural than head-mounted and smart-phone-like baselines (Konrad et al., 2024). G-VOILA interfaces with a textual LLM, using gaze-generated saliency maps for object detection. Derived object information is spliced into the query, increasing robustness against ambiguity and increasing participants’ confidence in the system (Wang et al., 2024). These prototypes show clear value from the inclusion of ET for point-in-time querying.
@@ 2.4. Our contributions
Our work further investigates ET’s role in contextual AI. We model the ET accuracy requirements to reliably detect objects, which is a prerequisite for agent understanding of the physical world. We also reveal added benefits when using historical gaze information to improve a contextual AI agent’s understanding of current context.
This work presents two key contributions: First, we estimate ET accuracy requirements needed for accurate gaze placement on physical objects. This investigation provides insights about the ET signal quality needed to robustly indicate user attention in the physical world. The usefulness of ET-informed contextual agents is limited not only by ET accuracy, however, but also by how information is conveyed to AI agents and how effectively the agents make use of this information. So secondly, we investigate the impact of including scanpath information as context for end-to-end contextual AI tasks. We supplement VLM queries with context sourced from the current and prior ET signals, augmenting agentic ability to understand user attention and current actions.
@@ 3. Methodology
To better understand ET’s role in future contextual AI systems, we first estimate the ET signal quality needed for accurate gaze-based selection of physical objects. Next, to measure the ability of AI agents to incorporate ET context, we experiment with end-to-end contextual AI tasks. We supplement VLM queries with contextual cues about current and prior ET signals, improving the model’s ability to understand user attention and current actions.
@@ 3.1. Eye tracking requirements for registering nearby objects
Contextual AI models will require knowledge of users’ real-world interests. ET is a primary signal to capture the user’s attention at a point in time. We hypothesize that the visual angle subtended by objects that users “look at” defines a lower bound on ET signal accuracy, and that such a system will require sufficient ET accuracy to consistently track the user’s point of focus. This lower bound is a prerequisite to then convey user attention to AI agents.
To investigate accuracy requirements, we analyze objects nearby in the user’s field of view (FOV) during natural tasks, relating the object size statistics to ET signal quality requirements. Individuals are far more likely to look at or interact with objects in the immediate vicinity (Ballendat et al., 2010), so we constrain our analysis to objects which are nearby candidates for interaction.
@@ 3.1.1. Dataset
For our analyses, we use a subset of the Aria Digital Twins (ADT) dataset (Pan et al., 2023). The ADT dataset contains egocentric recordings of daily-life tasks. We analyze single-participant recordings in the fully furnished apartment scene (93 recordings), totaling $\sim$3 hours of footage. Designed to model real-world household scenarios, these recordings span the following tasks: decorating, cooking, working, cleaning, and object examination.
In addition to the ET signal (median error $=1.5\degree$) provided by Project Aria glasses (Engel et al., 2023), the ADT dataset contains additional ground-truth information about physical objects in the scene. With motion capture markers and digitally twinned objects, each object’s position, orientation, bounding box, and segmentation area is computed and available at all points in time, with median tracking error of 5mm. This ground-truth information uniquely enables us to analyze object visual statistics at a fine scale, detect human-object interactions, and accurately place gaze on objects. 396 household objects are present across the dataset, with varying presence across different tasks and recordings.
@@ 3.1.2. Object visual size in relation to eye tracking error
Figure 1. Illustration of eye tracking spatial error and object visual size measurements. As an average case measurement, object segmentation region can be considered a circular region, with the radius reflecting the eye tracking accuracy requirement (thin bars). Alternatively, $1/2$ minor axis span $L_{min}$ (thick bars) represents a lower bound which considers non-uniform objects.
There is a relationship between an eye tracker’s spatial accuracy and the size of physical objects which can be reliably tracked, as is the case with virtual objects (Feit et al., 2017). At the same time, an objects’ visual size occupied within an image capture is inversely proportional to its distance from the user. Object visual size is the visual angle spanned by the object relative to the user’s FOV. To jointly account for distance and object physical size, we use the object’s visual size as a metric to predict ET accuracy needs.
@@ 3.1.3. Protocol
Using object segmentation information from the ADT dataset, we approximate the ET requirements in natural tasks / household environments by measuring the overall distribution of object visual spans. To better inform various contextual AI applications, we specify the ET requirements across different interaction spaces, namely:
@@
    Interacted objects: all objects being physically interacted with (held, pressed, pushed, etc.) by the participants, with an additional start / stop padding of 1 second.
@@ 3.2. Contextualizing vision-language model queries with eye tracking information
To showcase the value that ET signals provide in contextual AI systems, we model experiments which reflect potential end-to-end contextual AI systems. In these experiments, we construct historical context from past fixations on physical objects, measuring the impact of including such context in VLM queries.
We use the Meta Llama 3.2 90B VLM [^3] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ (Grattafiori et al., 2024) as our contextual AI agent. We supply context with an egocentric image and additional prompting. In both experiments, the agent is constrained via JSON response to output one currently visible object, as we are operating under the pretense that in a full system, an object recognition / scene understanding model would be available. For each prior fixation point being supplied as context, we inform the model which object was being gazed at. Note that this VLM is not fine-tuned, and that a model tuned for egocentric image understanding and / or for a specific task would see improved results. Yet, these experiments show the added value when incorporating ET contextual information.
@@
Figure 2. An example query to the VLM for E1. The additional context from fixation history is highlighted in red; this context is adjusted, evaluating the model’s ability to leverage various amounts of gaze context.
In this experiment, we pose the question “what am I looking at?” This experiment serves as a benchmark for the effect that prior eye gaze context serves in improving image understanding. We first detect and localize fixations on objects (using velocity-thresholding at 100°$\degree$ per second (Salvucci and Goldberg, 2000), and only considering fixations $\geq$ 150 ms), then perform uniform random sampling across each clip in the ADT dataset to analyze 919 image frames which each contain at least 10 prior fixations. At each sample, we make multiple queries, varying the number of prior fixations supplied as context to the VLM, between 0 - 10. An example prompt set can be seen in Figure 2.
@@
We query the VLM “what am I going to interact with?” while again varying historical gaze information from prior fixations. In this experiment, we supply the image from a current fixation in which a physical interaction is guaranteed to occur within the next second (N=237). The vast majority of interactions in ADT are grasping, pushing, and pulling with hands. This task models the use of ET as supporting context for user action understanding and prediction.
@@ 4.1. Eye tracking accuracy requirements
Figure 3. Eye tracking accuracy requirements in order to place gaze accurately onto the objects present in the ADT dataset, where users performed household actions in an indoor environment. Near-field and mid-field measurements consider all objects within the user’s field of view, where interacted objects are being actively manipulated by the user, and fixated objects consider those being directly gazed upon.
Objects which form a smaller image on the camera sensor require more accurate ET. To estimate ET accuracy requirements, we present the entire distribution of object visual sizes recorded in camera projection space. To place gaze on an object of average (projected) size 50% of the time, we measure at the distribution’s 50% mark; for more stringent coverage, a target further along the distribution may be chosen. The distributions for each interaction scenario are seen in Figure 3.
The near-field ($\leq$ 1 meter) and mid-field (1 - 2 meters) measurements reflect the visual FOV occupied by every object in the environment within the distance threshold. Alternatively, the interacted measurement considers only the objects being manually interacted with by the user, including picking up, pushing, pressing, etc. Fixation measurements consider the objects within 2 meters at the time of fixation. Considering ADT’s natural scenarios and environment, the interacted and fixation categories reflect the distribution of objects likely to be of interest during daily tasks.
Recall that we use angular radius measurements to represent average case ET requirements, and minor axis span to represent a more conservative measurement. The accuracy of wearable ET devices is known to suffer in natural, dynamic conditions (Onkhar et al., 2023), yet recent devices remain quite accurate in unconstrained settings [^4] https://www.tobii.com/products/eye-trackers/wearables/tobii-pro-glasses-3555https://pupil-labs.com/products/neon. Assuming a device with $\leq 3\degree$ accuracy during daily wear, our results indicate that the majority of fixated objects (radius average=$4.07\degree$; minor axis=$3.12\degree$), the majority of objects in the near-field (radius=$5.88\degree$; minor axis=$4.69\degree$), and nearly all interacted objects (radius=$10.81\degree$; minor axis=$9.10\degree$) are reliable for placing gaze on the correct object. Conversely, objects in the mid-field (radius=3.3°3.3°$\degree3$; minor axis=$2.54\degree$) will be somewhat unreliable at this signal quality, where half of objects are not able to be detected.

### v2

@@ Abstract.
Advanced multimodal AI agents can now collaborate with users to solve challenges in the world. Yet, these emerging contextual AI systems rely on explicit communication channels between the user and system. We hypothesize that implicit communication of the user’s interests and intent would reduce friction and improve user experience in contextual AI. In this work, we explore the potential of wearable eye tracking to convey user attention to the agents. We measure the eye tracking signal quality requirements to effectively map gaze traces to physical objects, then conduct experiments to provide visual scanpath history as additional context when querying multimodal agents. Our results show that eye tracking provides high value as a user attention signal and can convey information about the user’s current task and interests to the agent.
@@ 1. Introduction
Artificial intelligence (AI) agents have become more connected with users in daily life (Wienrich and Latoschik, 2021), especially by observing context about the user’s prior actions or current world state (Zhang et al., 2024a). New innovations, such as vision-language models (VLMs) (Li et al., 2024) and machine perception devices (Engel et al., 2023), pave the way towards contextual AI agents: agents which “see” the nearby physical world and collect contextual cues to better assist users. Yet, current models interpret context differently than humans, so often misinterpret context, conflicting with user intent. If implicit information about the user’s state could be reliably supplied to the agent, the user and agent’s intent could be better aligned.
Eye gaze has been hypothesized as a valuable signal for conveying intent to agents (Ajanki et al., 2010; Büschel et al., 2018; Burlingham et al., 2024b; Zhang et al., 2024b). Gaze conveys information about objects users are interested in, cognitive load, the current action, etc. (Mahanama et al., 2022), all of which could improve models’ understanding. While eye tracking (ET) is a common input in extended reality (XR) systems (Plopski et al., 2022), the use of ET in human-agent interactions has only been lightly explored (Sendhilnathan et al., 2024).
ET signals could convey to an agent what the user is or has been interested in. Wearable eye trackers are limited in accuracy due to multiple factors (system error, slippage, individual user differences, etc.) (Ehinger et al., 2019), constraining whether objects could be reliably identified. If an object’s visual size relative to the ET sensor is too small, it could be unreliable to detect.
We present an analysis of the requirements and benefits of ET in wearable contextual AI. Using a dataset of egocentric recordings taken during daily household tasks (Pan et al., 2023), we estimate the expected ET accuracy thresholds for detecting physical objects in different contexts. We then conduct a number of experiments where contextual information from ET is appended to VLM queries. These experiments reinforcing ET’s value in this space, greatly improving the model’s ability to perceive user attention and current actions.
@@ 2.1. Eye tracking for selection in extended reality
In recent years, eye gaze has gained popularity as a key signal for HCI in XR systems (Plopski et al., 2022). Eye gaze has comparable usability to controllers (Luro and Sundstedt, 2019; Zhang et al., 2019; Fernandes et al., 2024) while freeing the hands for other tasks and being preferable to users (Piening et al., 2021). While ET is prone to a midas touch fallacy, where false selections are made during ambient fixations (Jacob, 1995), novel HCI methodologies (Khamis et al., 2018) overcome this and make ET an ideal signal for interface navigation.
Our work explores ET’s ability to continuously and implicitly convey a user’s attention and intent (Sendhilnathan et al., 2024). This has only lightly been explored with AI agents, though ET has facilitated automatic contextual displays in the past (Toyama et al., 2012).
@@ 2.2. Eye gaze encodes scene understanding
Eye movements reflect a viewer’s internal processing of a scene (Yarbus, 1967), giving insights to cognitive state and attention as one interprets new visual stimuli (Langton et al., 2000; Eckstein et al., 2017). Sequences of gaze fixations (i.e., scanpaths) encodes contextual cues as to future objects of interest (Itti and Koch, 2001; Burlingham et al., 2024a); a number of works have leveraged scanpath history for short-term gaze prediction / anticipation (Huang et al., 2018; David-John et al., 2021; Hu et al., 2021; D’Amelio et al., 2024). Burlingham et al. found temporal dependencies in scanpaths lasting for 4-5 fixations on average, with high variance across task types (Burlingham et al., 2024b). Contextual AI models may be able to leverage the rich, multiscale structure of scanpaths (analogous to how large language models (LLMs) leverage multiscale structures in language), enabling implicit inferences about user intent from scanpath history.
Insights about the cognitive encoding of nearby objects can inform our expectations for eye movements in contextual AI (Tatler et al., 2011). For example, humans tend to look at a coffee mug just before grasping, to encode the location of the handle so that it can be successfully grasped. Object locations are encoded via an egocentric reference to the user, and greater affordances are given to nearby interactable objects (Costantini et al., 2010; Tatler et al., 2011). The visual system elicits responses to reachable 3D objects, even when there is no intent to interact (Iachini et al., 2014, 2023). So, by analyzing the eye gaze fixations on nearby objects, we are able to register a list of possible future interactions.
@@ 2.3. Eye tracking in contextual AI
Information from the physical world could greatly improve user interactions with AI agents (Zhang et al., 2024a). Emerging products, such as the Ray-Ban Meta [^1] https://www.meta.com/smart-glasses and Google’s Ask Photos [^2] https://blog.google/products/photos/ask-photos-google-io-2024/, use image context to improve user interaction. But, given a full image, the human’s intent may not align with the salient features detected by the agent. Eye gaze could be fundamental for narrowing relevant context observed by contextual agents (Ajanki et al., 2010; Büschel et al., 2018), enhancing understanding and avoiding hallucination (Cui et al., 2023; Leng et al., 2024).
Some wearable contextual AI prototypes integrating eye tracking have been proposed (Zhang et al., 2024b). The GazeGPT system projects 2D gaze onto an image capture, cropping image contents before interfacing with a VLM (Konrad et al., 2024). They show gaze-based querying to be faster, more accurate, and more natural than head-mounted and smart-phone-like baselines. G-VOILA interfaces with a textual LLM, using gaze-generated saliency maps for prompt adjustment (Wang et al., 2024). Derived object information is spliced into the query, increasing robustness against ambiguity and increasing participants’ confidence in the system. These prototypes show clear value from the inclusion of ET for point-in-time querying.
@@ 2.4. Our contributions
Our work further investigates ET’s role in contextual AI. We model the ET accuracy requirements to reliably detect objects, which is a prerequisite for detecting user attention in the physical world. We also explore the added benefits of historical gaze scanpaths to improve a contextual AI agent’s understanding of current context.
This work presents two key contributions: First, we estimate ET accuracy requirements needed for accurate gaze placement on physical objects, to determine ET accuracy requirements for future systems. Secondly, we explore how information is conveyed to AI agents, investigating temporal dependencies rather than point-in-time gaze. We include scanpath information as context for end-to-end contextual AI tasks, and augment VLM queries with this context. This augments agentic ability to understand user attention and current actions.
@@ 3. Methodology
To better understand ET’s role in future contextual AI systems, we first estimate the ET signal quality needed for accurate gaze-based selection of physical objects. Next, to measure the ability of AI agents to incorporate ET context, we experiment with end-to-end contextual AI tasks. We inject contextual cues derived from the user’s eye gaze into VLM queries, improving the model’s ability to understand user attention and current actions.
@@ 3.1. Eye tracking requirements for registering nearby objects
Contextual AI models that collaborate with users would benefit from knowledge of users’ real-world interests. ET is a promising signal to capture the user’s attention, both at one point in time (Konrad et al., 2024), or continuously to build historical context. We hypothesize that the visual angle subtended by objects that users “look at” defines a lower bound on ET signal accuracy. By measuring the visual angle expressed by nearby objects, we can approximate the ET accuracy required to consistently track a user’s point of focus.
To investigate accuracy requirements, we analyze objects that are nearby in the user’s field of view (FOV) during daily household tasks, then relate the object size statistics to ET signal quality requirements. Because we expect humans and AI agents to collaborate in daily life, it is important for the ET system to achieve an accuracy which allows consistent, accurate attention modeling in a broad set of scenarios (Feit et al., 2017). Individuals are far more likely to look at or interact with objects in the immediate vicinity (Ballendat et al., 2010), so we constrain this analysis to objects which are nearby candidates for interaction.
@@ 3.1.1. Dataset
For our analyses, we use a subset of the Aria Digital Twins (ADT) dataset (Pan et al., 2023). The ADT dataset contains egocentric recordings of daily-life tasks in indoor environments. We analyze single-participant recordings in the fully furnished apartment scene (93 recordings), totaling $\sim$3 hours of footage. Designed to model real-world household scenarios, these recordings span the following tasks: decorating, cooking, working, cleaning, and object examination.
In addition to the ET signal (median error $=1.5\degree$) provided by Project Aria glasses (Engel et al., 2023), the ADT dataset contains ground-truth information about physical objects in the scene. Each object’s position, orientation, bounding box, and segmentation region is tracked and available throughout a recording, with median tracking error of $5mm$. This ground-truth data uniquely enables us to analyze object visual statistics at a fine scale, detect human-object interactions, and accurately place gaze on objects. 396 household objects are tracked in the dataset, with varying presence across different tasks and recordings.
@@ 3.1.2. Object visual size in relation to eye tracking error
Figure 1. Illustration of eye tracking spatial error and object visual size measurements. As an average case measurement, object segmentation area can be mapped to a circular region, with the radius reflecting the eye tracking accuracy requirement (thin bars). Alternatively, $1/2$ minor axis span $L_{min}$ (thick bars) is a tighter bound for measuring non-uniform objects.
ET spatial error is the measured bias between the ground truth and estimated gaze positions. This error persists following techniques such as ET calibration and fixation detection. We measure spatial error as an angular offset between the user’s true gaze point and computed value. We wish to approximate the influence of ET spatial error in placing fixations on object, to better inform how reliably an ET system could convey a user’s attention on world objects.
@@ 3.1.3. Protocol
We approximate the ET requirements for daily use by observing the household tasks being performed in the ADT dataset (Pan et al., 2023). We measure the distributions of object visual sizes for $err_{{}_{ET}}$ and $err_{{}_{ET}\ low}$. To better inform various contextual AI applications, we specify the ET requirements across different interaction spaces, namely:
@@
    Interacted objects: all objects being physically interacted with (held, pressed, pushed, etc.) by the participants, with a start / stop padding of 1 second for the interaction.
@@
Figure 2. An example query to the VLM for E1. The additional context from fixation history is highlighted in red; We vary the amount of context given to the model to measure how this context influences model understanding.
@@ 3.2. Contextualizing vision-language model queries with eye tracking information
To begin to explore the value that ET signals can provide in contextual AI systems, we model experiments which reflect potential end-to-end contextual AI systems. In these experiments, we build up historical context by creating timelines of past fixations on physical objects. We include this context in VLM queries and measure the positive impacts on model understanding. Our baseline comparison is a VLM query which uses only the egocentric image as added context (similar to a Ray-Ban Meta or Google Ask Photos query).
We use the Meta Llama 3.2 90B VLM [^3] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ (Grattafiori et al., 2024) as our contextual AI agent. We add context to queries with an egocentric image and additional prompting. In both experiments, the agent is constrained via JSON to respond with one option from all currently visible objects, as we are operating under the pretense that in a full system, an object recognition / scene understanding model would be available. Note that this VLM is not fine-tuned, and that a model tuned for egocentric image understanding and / or for a specific task would likely see improved results. Yet, these experiments indicate the added value when incorporating ET contextual information.
@@
In this experiment, we pose the question “what am I looking at?” This experiment serves as a benchmark for the effect that prior eye gaze context serves in improving image understanding. We first detect and localize fixations on objects [^4] We use a velocity-thresholding algorithm at 100°$\degree$ per second (Salvucci and Goldberg, 2000), to account for the relatively low sampling rate of Project Aria glasses (30Hz) (Engel et al., 2023). We only consider fixations $\geq$ 150 ms for analysis., then perform uniform random sampling across each recording in the ADT dataset to analyze 919 image frames which each contain at least 10 prior fixations. At each sample, we make multiple queries, varying the number of prior fixations supplied as context to the VLM, between 0 - 10. An example prompt set can be seen in Figure 2.
@@
We query the VLM “what am I going to interact with?” while again varying the amount of gaze context. In this experiment, we supply the image from a current fixation, where a physical interaction is guaranteed to occur within the next second and at least 10 prior fixations exist (237 distinct image frames). The vast majority of physical interactions in ADT are grasping, pushing, and pulling with hands. This task models the use of ET as supporting context for user action understanding and prediction.
@@ 4.1. Eye tracking accuracy requirements
Figure 3. Eye tracking accuracy requirements to place gaze accurately within in the ADT dataset, where users performed household actions in an indoor environment (Pan et al., 2023) Near-field and mid-field measurements consider all objects present in the user’s field of view, where interacted objects are being actively manipulated by the user, and fixated objects consider those being directly gazed at.
The near-field ($\leq$ 1 meter) and mid-field (1 - 2 meters) measurements reflect the visual FOV occupied by every object in the environment within the distance threshold. Alternatively, the interacted measurement only aggregates objects when being manually interacted with, including picking up, pushing, pressing, etc. Fixation measurements consider the objects within 2 meters at the time of fixation. Considering ADT’s household scenarios, the interacted and fixation categories reflect the distribution of objects likely to be of interest during daily tasks.
To estimate ET accuracy requirements, we computed the entire distribution of object visual sizes recorded in camera projection space. To place gaze on an object of average (projected) size 50% of the time, we measure at the distribution’s 50% mark. These measurements can help to inform system design; while 50% reliability may be useful supporting context in a broader contextual AI agent, a system relying heavily on ET may aim for higher coverage. The distributions for each interaction scenario are seen in Figure 3.
Recall that we use angular radius to represent the average case ET requirements, and minor axis span as a more conservative measurement. The accuracy of wearable ET devices is known to suffer in dynamic conditions (Onkhar et al., 2023), yet recent devices remain quite accurate in unconstrained settings [^5] https://www.tobii.com/products/eye-trackers/wearables/tobii-pro-glasses-3666https://pupil-labs.com/products/neon. Assuming a device with $\leq 3\degree$ accuracy during daily wear, our results indicate that the majority of fixated objects (radius average=$4.07\degree$; minor axis=$3.12\degree$), the majority of objects in the near-field (radius=$5.88\degree$; minor axis=$4.69\degree$), and nearly all interacted objects (radius=$10.81\degree$; minor axis=$9.10\degree$) are reliable for placing gaze on the correct object. Conversely, objects in the mid-field (radius=3.3°3.3°$\degree3$; minor axis=$2.54\degree$) will be somewhat unreliable at this signal quality, where roughly half of objects are not able to be detected.